OpenStudy (anonymous):

A mechanical engineer bought 24 boxes of screws for P2200. There were three types of screws bough. Screw A cost P300 per box, screw B costs P150 per box and screw C costs P50 per box. How many boxes of each screw did he buy?

OpenStudy (ajprincess):

Let the number of boxes of screw A be x , the number of boxes of screw B be y and the number of boxes of screw C be z. Then x+y+z=24 300x+150y+50z=2200 30x+15y+5z=220 6x+3y+z=44. I feel like some data is missing.

OpenStudy (anonymous):

Yeah, you'd expect there be a third equation... There is some data that can help us narrow things down though. For instance: \(x, y, z \in \mathbb{N} \)

OpenStudy (anonymous):

That would exclude 0.

OpenStudy (anonymous):

{x = 2, y = 5, z = 17} Solved @ajprincess' first two equations in 1/1000 sec using Mathematica's Constrained Optimization function, Maximize. Refer to the attachment.